The length of a rectangle is 2 meters more than its width. If the area is 35 square meters, what are the length and width?

Answer by ewatrrr(24785)


Hi
length of a rectangle is 2 meters more than its width.
Let x and (x+2) represent the width and length respectively
Question states***
x(x+2) = 35 m^2
Solving for x
x^2 + 2x - 35 = 0
factoring
(x+7)(x-5)= 0 Note:SUM of the inner product(7x) and the outer product(-5x) = 2x
(x+7)=0 |tossing out negative solution for length
(x-5)= 0 x = 5m, the width. The length is 7m. (5m + 2m)
CHECKING our Answer***
5m*7m = 35 m^2